OpenStudy (anonymous):

9x - 3(6 – 5x) = x – 2(3x + 9)

OpenStudy (anonymous):

Isolate your x

OpenStudy (anonymous):

Do your brackets first though.

OpenStudy (anonymous):

how do i do that am i subtracting it or combining like terms

OpenStudy (anonymous):

Let's do it step by step.

OpenStudy (anonymous):

1rst, undo your brackets. You know how?

OpenStudy (anonymous):

3(6-5x) 3 x 6 = ? 3 x -5x = ?

OpenStudy (anonymous):

9x - 3(6 – 5x) = x – 2(3x + 9) 9x-18-15x=x-6x-18......like this

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

Although the last one is +18

OpenStudy (anonymous):

From here, you can bring your X's on one side and the numbers on the other.

OpenStudy (anonymous):

like combine the like terms

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

how do i get rid of the equal sign

OpenStudy (anonymous):

You don't get rid of it, you keep it.

OpenStudy (anonymous):

You want to end up with x =?

OpenStudy (anonymous):

can you write it out for im a little lost on how to show my work wen combining the terms

OpenStudy (anonymous):

so i would do 9x-15x+x

OpenStudy (anonymous):

Ok look, you had your equation: 9x-18-15x=x-6x+18 This becomes -6x-18 = -5x +18 Next, you bring the -5x over so it becomes positive. -x-18 = 18 Bring the -18 over -x = 36 x=-36

OpenStudy (anonymous):

- times 9 =-18 its not postive

OpenStudy (anonymous):

-2 times 9 =-18

OpenStudy (anonymous):

It's positive. The negative is between the two sets and is not with the number itself.

OpenStudy (anonymous):

If it would have been x + - 2(3x + 9) or x – -2(3x + 9) , then yes

OpenStudy (anonymous):

still dont see how you got 36

OpenStudy (anonymous):

-36

OpenStudy (anonymous):

I'll explain it in details then.

OpenStudy (anonymous):

i put the whole problem in on caluclator and it gave me 0 for x

OpenStudy (anonymous):

EXAMPLE: 2x + 3 = x -2 (2x-x) +3 = (x-x) -2 x+3 = -2 x+3-3 = -2-3 x=-5 Same thing for your equation! 9x - 3(6 – 5x) = x – 2(3x + 9) Multiply the brackets! 3(6-5x) = 3 x 6 = 18 3 x -5x = -15x 2(3x+9)= 2 x 3x = 6x 2 x 9 = 18 Now you have your new equation!!!!!! 9x - 18-15x = x - 6x + 18 Now collect like-terms! 9x-15x-18 = x-6x +18 -6x-18 = -5x +18 (-6x+5x)-18 = (-5x+5x) +18 -x-18 = 18 -x-18+18 = 18+18 -x = 36 --x=-36 x=-36

OpenStudy (anonymous):

I hope this helped because I can't explain it any other way.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you help me with another one

OpenStudy (anonymous):

Sure, open a new question up

OpenStudy (anonymous):

k

OpenStudy (anonymous):

You have to close this question first though.